### 08 Jan 2024 v1

@@ 4.2 Computational efficiency
GrainGNN achieves a substantial reduction in the storage of variables and the number of required time steps, which leads to significant speedups over phase-field simulations. For the training domain size, the required number of phase field variables is about 500M despite using the active parameter tracking and moving-domain algorithms, while the number of features of GrainGNN is about 3.2K. The network has about 2.4M parameters. Thus, GrainGNN requires $10^{2}-10^{5}$ times less storage than the phase field solver. The computation efficiency of our phase field solver and GrainGNN is listed in Table 2. Our phase field solver is optimized with GPUs [53]. For the $G_{z}$ and $R_{z}$ values investigated in this paper, the time cost per phase field simulation is 300--3000 seconds on one A100 GPU, while the time cost of GrainGNN is 0.2--3 seconds. GrainGNN on a single CPUachieves 150$\times$–2000$\times$ speedup over our phase field code on a single A100 GPU.
@@
Table 2: Computational efficiency of GrainGNN for different $\boldsymbol{p}$ and domain size $L$. Time for solving phase field equations was measured on one NVIDIA A100 GPU with 40 GB memory. The domain size is $(40\mu m,40\mu m,50\mu m)$. GrainGNN inference time was measured on a single MD EPYC 7763 CPU core. $a_{x}\times a_{y}\times a_{z}$ represents the GrainGNN inference time for a domain size of $(a_{x}L^{0}_{x},a_{y}L^{0}_{y},a_{z}L^{0}_{z})$. Every measured time is averaged across 10 initial realizations of grains.
@@ 4.3 Extensions and limitations
There are several planned extensions to the current framework of GrainGNN. One is generalizing the computational domain to no-flux boundary conditions and non-rectangular geometries. For no-flux boundary conditions, we consider padding domain boundaries with halo grains that have mirrored properties (e.g., orientation) with respect to the grains on the boundary. The number of padding grains depends on the domain and physical parameters. For domain geometry, we plan to follow the rectangular-to-curvilinear domain mapping strategy presented in GrainNN [17]. The idea is to find geometric coefficients that map a curved surface to a 2D plane. We run GrainGNN in a rectangular domain and use the geometric coefficients to scale the network outputs for example the junction displacements. The output structure is mapped back to the original geometry. A second extension is to improve the representations of grain boundaries. We neglect the in-plane curvature of the grain boundaries, which can be significant for high $R_{z}$ values. Another goal is to use experimental design and active learning [62, 63] methods for sampling physical parameters used for training. Currently, we use a uniform grid of $G_{z}$ and $R_{z}$ to generate training data. The MR plot in Fig. 9 indicates that more data should be drawn from the region with high $R_{z}$ and low $G_{z}$ to reduce the pointwise error. To make the data generation computationally trackable for higher dimensional parameter space, an efficient adaptive sampling algorithm needs to be developed. From the network architecture perspective, the current one-to-one LSTM prediction can be extended to sequence-to-sequence prediction to improve accuracy as we did in [17]. However is unclear how to handle topological changes in sequence-to-sequence configurations.
@@
Limitations. The two major limitations of GrainGNN are dealing with complex meltpool geometries and ignoring grain nucleation. We discussed geometry in the previous paragraph. Grain nucleation introduces new grains to the system thus adding vertices and edges to the graph. The modifications of the graph are two kinds depending on the nucleation density. For dilute nucleation, we can still utilize GrainGNN. We will introduce a new operation $O_{N}$ that adds vertices and edges, which will be the inverse operation of $O_{G}$. For dense nucleation, the equiaxed growth dominates the grain formation. This evolution is substantially different enough from epitaxial growth that a third network (beyond $\mathcal{R}$ and $\mathcal{C}$) will have to be added to predict the graph generation during grain nucleation. A third more fundamental limitation is detecting failure cases, in short some kind of a posteriori error estimates for network predictions. GrainGNN doesn’t have any performance guarantees other than the empirical evaluation we discussed. A forth limitation is that the inference phase of GrainGNN runs only CPUs, and thus potential speedup are still possible. We are currently working to address these challenges.
@@ 5 Conclusions
We presented a surrogate for microstructure evolution in 3D epitaxial grain formation. We introduced a heterogeneous graph model and hand-crafted graph features that combined achieve a significant spatiotemporal compression of grain microstructure. We proposed an image-to-graph method to extract graphs from phase field data. We modeled microstructure formation with graph-to-graph evolution, where we decomposed the evolution into feature changes and topological events. GrainGNN is implemented with an LSTM-based regressor and classifier to predict the features and events. The LSTMs include graph transformers and can scale to a larger number of grains than those used for training. We also proposed graph and microstructure reconstruction methods to address topological changes. GrainGNN is trained with phase field data in a wide range of $G$ and $R$ values; in AM process conditions. It can predict both quantities of interest and pointwise accurate microstructure for unseen process parameters and initial grain configurations. We also showed that GrainGNN generalizes to domain size, number of grains, and initial grain parameters. GrainGNN is orders of magnitude faster than high fidelity simulations and scalable to a large number of grains.
@@ Appendix A Phase field solver
symbol meaning value [6, 4]
@@
$L/C_{p}$ latent heat/heat capacity(K) 229
$\lambda$ thermal coupling constant 58.3
@@
$D$ heat diffusion coefficient ($m^{2}/s$) $3.6\times 10^{-6}$
@@
Fig. A.1: Convergence test for a phase field simulation. $G_{z}=10K/\mu m$ $R_{z}=2m/s$. The number of grains is 463. The simulation parameters used for generating the training data (blue line) are interface width $W_{0}=1\mu m$, mesh size $dx=0.8W_{0}$, moving-domain height $L_{z}=8\mu m$, and five phase field variables per grid point. The yellow and green lines show the convergence of the discretizations by halving $W_{0}$ or $dx$. The red line shows the convergence of the moving-domain algorithm by doubling the moving-domain width. The convergence of the blue and purple lines indicates the negligible difference between storing five phase fields and ten phase fields.
inline]add discussion about KS and how to interpret its values.
inline]Code description and gihub repo etc must be available.

### 31 Jan 2024 v2

@@ 4.2 Computational efficiency
GrainGNN achieves a substantial reduction in the storage of variables and the number of required time steps, which leads to significant speedups over phase field simulations. For the training domain size, the required number of phase field variables is about 500M despite using the active parameter tracking and moving-domain algorithms, while the number of features of GrainGNN is about 3.2K. The network has about 2.4M parameters. Thus, GrainGNN requires $10^{2}-10^{5}$ times less storage than the phase field solver. The computation efficiency of our phase field solver and GrainGNN is listed in Table 3. Our phase field solver is optimized with GPUs [53]. For the $G_{z}$ and $R_{z}$ values investigated in this paper, the time cost per phase field simulation is 300--3000 seconds on one A100 GPU, while the time cost of GrainGNN is 0.2--3 seconds. GrainGNN on a single CPU achieves 150$\times$–2000$\times$ speedup over our phase field code.
@@
Table 3: Computational efficiency of GrainGNN for different $\boldsymbol{p}$ and domain size $L$. Time for solving phase field equations was measured on one NVIDIA A100 GPU with 40 GB memory. The domain size is $(40\mu m,40\mu m,50\mu m)$. GrainGNN inference time was measured on a single AMD EPYC 7763 CPU core. $a_{x}\times a_{y}\times a_{z}$ represents the GrainGNN inference time for a domain size of $(a_{x}L^{0}_{x},a_{y}L^{0}_{y},a_{z}L^{0}_{z})$. Every measured time is averaged across 10 initial realizations of grains.
@@ 4.3 Extensions and limitations
There are several planned extensions to the current framework of GrainGNN. One is generalizing the computational domain to no-flux boundary conditions and non-rectangular geometries. For no-flux boundary conditions, we consider padding domain boundaries with halo grains that have mirrored properties (e.g., orientation) with respect to the grains on the boundary. The number of padding grains depends on the domain and physical parameters. For domain geometry, we plan to follow the rectangular-to-curvilinear domain mapping strategy presented in GrainNN [17]. The idea is to find geometric coefficients that map a curved surface to a 2D plane. We run GrainGNN in a rectangular domain and use the geometric coefficients to scale the network outputs for example the junction displacements. The output structure is mapped back to the original geometry. A second extension is to improve the representations of grain boundaries. We neglect the in-plane curvature of the grain boundaries, which can be significant for high $R_{z}$ values. Another goal is to use experimental design and active learning [63, 64] methods for sampling physical parameters used for training. Currently, we use a uniform grid of $G_{z}$ and $R_{z}$ to generate training data. The MR plot in Fig. 10 indicates that more data should be drawn from the region with high $R_{z}$ and low $G_{z}$ to reduce the pointwise error. To make the data generation computationally trackable for higher dimensional parameter space, an efficient adaptive sampling algorithm needs to be developed. From the network architecture perspective, the current one-to-one LSTM prediction can be extended to sequence-to-sequence prediction to improve accuracy as we did in [17]. However it is unclear how to handle topological changes in sequence-to-sequence configurations.
@@
Limitations. The two major limitations of GrainGNN are dealing with complex meltpool geometries and ignoring grain nucleation. We discussed geometry in the previous paragraph. Grain nucleation introduces new grains to the system thus adding vertices and edges to the graph. The modifications of the graph are two kinds depending on the nucleation density. For dilute nucleation, we can still utilize GrainGNN. We will introduce a new operation $O_{N}$ that adds vertices and edges, which will be the inverse operation of $O_{G}$. For dense nucleation, the equiaxed growth dominates the grain formation. This evolution is substantially different enough from epitaxial growth that a third network (beyond $\mathcal{R}$ and $\mathcal{C}$) will have to be added to predict the graph generation during grain nucleation. A third more fundamental limitation is detecting failure cases, in short some kind of a posteriori error estimates for network predictions. GrainGNN doesn’t have any performance guarantees other than the empirical evaluation we discussed. A fourth limitation is that the inference phase of GrainGNN runs only CPUs, and thus potential speedups are still possible. We are currently working to address these challenges.
@@ 5 Conclusions
We presented a surrogate for microstructure evolution in 3D epitaxial grain formation. We introduced a heterogeneous graph model and hand-crafted graph features that combined to achieve a significant spatiotemporal compression of grain microstructure. We proposed an image-to-graph method to extract graphs from phase field data. We modeled microstructure formation with graph-to-graph evolution, where we decomposed the evolution into feature changes and topological events. GrainGNN is implemented with an LSTM-based regressor and classifier to predict the features and events. The LSTMs include graph transformers and can scale to a larger number of grains than those used for training. We also proposed graph and microstructure reconstruction methods to address topological changes. GrainGNN is trained with phase field data in a wide range of $G$ and $R$ values in AM process conditions. It can predict both quantities of interest and pointwise accurate microstructure for unseen process parameters and initial grain configurations. We also showed that GrainGNN generalizes to domain size, number of grains, and initial grain parameters. GrainGNN is orders of magnitude faster than high fidelity simulations and scalable to a large number of grains.
@@
Acknowledgments
This material is based upon work partially supported by NSF award OAC 2204226 and by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Applied Mathematics program, Mathematical Multifaceted Integrated Capability Centers (MMICCS) program, under award number DE-SC0023171. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily reflect the views of the DOE, and NSF. Computing time on the Texas Advanced Computing Centers Stampede system was provided by an allocation from TACC and the NSF.
@@ Appendix A Phase field solver
The material parameters of stainless steel 316L are listed in Table A.1. Given values of $\Gamma$, $L_{f}/C_{p}$, and $\mu_{k}$, phase field parameters $\lambda$ and $\tau_{0}$ can be derived using asymptotic analysis [56, 53]:
$\frac{\Gamma C_{p}}{L_{f}}\equiv d_{0}=a_{1}\frac{W_{0}}{\lambda},$ (A.1)
$\frac{C_{p}}{\mu_{k}L_{f}}\equiv\beta_{0}=a_{1}\frac{\tau_{0}}{\lambda W_{0}}-a_{1}a_{2}\frac{W_{0}}{D_{h}},$ (A.2)
where $a_{1}=5\sqrt{2}/8$, $a_{2}=47/75$; $d_{0}$ is the thermal capillarity length; $\beta_{0}$ is the kinetic coefficient; $D_{h}$ is the heat diffusion coefficient. The obstacle parameter $\omega$ in Eq. 2.1a is set to $\lambda u/I$, where $u$ is the nondimensional undercooling and $I$ is a constant, here we choose $I=1/12$ [3]. The anisotropy of $\tau_{\alpha}$ requires the spatial derivatives of $\phi$ in the rotated coordinates with angle $\boldsymbol{\theta}$. Let ${\theta}_{z}$ be the angle between the $z$-axis and the preferred growth direction ¡100¿ and ${\theta}_{x}$ be the angle between the $x$-axis and the projection of ¡100¿ direction on the $x$-$y$ plane. The derivatives appearing in Eq. 2.2b should be replaced with:
$\left[\begin{array}[]{c}{\nabla}_{x}\phi^{\prime}\\ {\nabla}_{y}\phi^{\prime}\\ {\nabla}_{z}\phi^{\prime}\end{array}\right]=\left[\begin{array}[]{ccc}\cos(\theta_{x})\cos(\theta_{z})&\sin(\theta_{x})\cos(\theta_{z})&-\sin(\theta_{z})\\ -\sin(\theta_{x})&\cos(\theta_{x})&0\\ \cos(\theta_{x})\sin(\theta_{z})&\sin(\theta_{x})\sin(\theta_{z})&\cos(\theta_{z})\end{array}\right]\cdot\left[\begin{array}[]{c}{\nabla}_{x}\phi\\ {\nabla}_{y}\phi\\ {\nabla}_{z}\phi\end{array}\right]$ (A.3)
Because the capillary anisotropy is much weaker than kinetic anisotropy in the rapid solidification regime, we set $\epsilon_{c}=0$.
The phase field solver is available at https://github.com/YigongQin/cuPF. The CUDA functions have been optimized. The most expensive right-hand side calculations have the performance of 270 GFlops on one Nvidia A100 GPU. We use CUDA-aware MPI which scales well on multiple GPUs and multiple nodes.
Fig. A.1 shows the convergence results for $dx$, $W_{0}$, the height of the moving domain, and the number of phase field variables per grid point used in the active parameter tracking algorithm. The training data is generated using the configuration: $W_{0}=0.1\mu m$, $dx=0.8W_{0}$, $L^{\prime}_{z}=8\mu m$, and five phase fields. We can see that the grain size distributions converge well when we halve the mesh size or increase the domain height or the number of phase fields.
@@
symbol meaning (units) value [6, 4]
$\Gamma$ Gibbs-Thompson coefficient (Km) $3.47\times 10^{-7}$
$L_{f}/C_{p}$ latent heat/heat capacity (K) 229
$\mu_{k}$ linear kinetic coefficient (m/s/K) 0.217
$\lambda$ thermal coupling constant 58.3
@@
$D_{h}$ heat diffusion coefficient ($m^{2}/s$) $3.6\times 10^{-6}$
@@
Fig. A.1: Convergence test for a phase field simulation. $G_{z}=10K/\mu m$ $R_{z}=2m/s$. The number of grains is 463. The simulation parameters used for generating the training data (blue line) are interface width $W_{0}=0.1\mu m$, mesh size $dx=0.8W_{0}$, moving-domain height $L^{\prime}_{z}=8\mu m$, and five phase field variables per grid point. The yellow and green lines show the convergence of the discretizations by halving $W_{0}$ or $dx$. The red line shows the convergence of the moving-domain algorithm by doubling the moving-domain width. The convergence of the blue and purple lines indicates the negligible difference between storing five phase fields and ten phase fields.
Appendix B Kolmogorov–Smirnov test
Kolmogorov-Smirnov test can be used to compare two samples and test whether the two samples could have come from the same distribution. If sample A has $m$ data points and sample B has $n$ data points, the null hypothesis that two samples coming from the same distribution is rejected when:
$\text{KS}>\sqrt{-\ln\left(\frac{\alpha}{2}\right)\cdot\frac{1+\frac{n}{m}}{2n}},$ (B.1)
where KS is the Kolmogorov-Smirnov statistic and $\alpha$ is the confidence level. For our setup, if a testing case has $n=120$ grains and we think GrainGNN and phase field results are the same with a 95% confidence level, the critical KS value is $\text{KS}_{c}=\sqrt{-\ln\left(\frac{0.95}{2}\right)/120}=0.079$. As discussed in Section 3.2, the KS values for in-distribution generalization are mostly in the range of 0.03--0.06, meaning that GrainGNN predicts the same grain size distribution as the phase field model. For out-of-distribution generalization we change the number of grains. $\text{KS}_{c}$ is 0.043 for $n=400$ and 0.022 for $n=1600$. In Fig. 10h we vary the initial orientation distribution with $n=400$. We can see $\text{KS}\approx\text{KS}_{c}$ for all $\theta_{0}$, which means we have 95% confidence GrainGNN and phase field have the same grain size distributions. In Fig. 10g when we decrease initial grain sizes, $\text{KS}>\text{KS}_{c}$ for $d_{0}<3\mu m$, which indicates a higher distribution mismatch for smaller grains.
Appendix C Code and data availability
GrainGNN codes are available at https://github.com/YigongQin/GrainGraphNN. Codes contain how to create a graph using the Voronoi diagram and how to extract a graph from grain microstructure images. The neural networks are developed based on the Pytorch Geometric Library. Training and testing of the networks with comparisons against the phase field results are provided. The GitHub repository also provides the trained models including the regressor and the classifier. The phase field data used for training can be reproduced with the CUDA codes at https://github.com/YigongQin/cuPF.